A. A mass of 5kg is on too of a building of height 10m. Calculate its potential energy.
B. The mass then falls at a speed of 8m/s in 4 secs, find its KE and momentum p. What is the force and impulse the mass transfers to the ground?
C. What is the work done by gravity if the mass falls a distance 2 meters. If the mass falls for 11 seconds, what is the energy expended by this force?
1
Expert's answer
2019-11-24T17:02:08-0500

A. The potential energy:


"E_P=mgh=5\cdot9.8\cdot10=490 \text{ J}."


B. The kinetic energy (when the velocity was 8 m/s):


"E_K=\frac{1}{2}mv^2=160\text{ J}."

The momentum:


"p=mv=40\text{ kg-m/s}."

The force:


"F=\frac{\Delta p}{\Delta t}=\frac{40}{4}=10\text{ N}."

The impulse;


"J=\Delta p=p-0=40\text{ kg-m/s}."

C. The work done by gravity:


"W=mgh=98\text{ J}."

The height required to fall within 11 s:


"h=gt^2\/2."

The energy expended by the force of gravity:


"E=Fh=mgh=mg^2t^2/2=29052.1\text{ J}."